OpenStudy (anonymous):

Can you work this problem out? I already know the right answer I just need to see how it was solved;; What is the relationship between the lines determined by the following two equations? 6x – 2y = 16 y = 3x – 8

OpenStudy (anonymous):

The answer is 'they are the same line' I just need to know how they got that answer.

OpenStudy (studygurl14):

Lines are the same exact line if they are identical. REarrange the first equation into the most simpllified slope-intercept form 6x - 2y = 16 -2y = -6x + 16 y = 3x - 8 Notice how this is now identical to the second equation.

OpenStudy (anonymous):

Yes. I see. Thank you

OpenStudy (studygurl14):

You're welcome.